Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Fri Jun 5 11:57:54 METDST 1998) Minor failures
No apparent problems until 8:32 UT. At this moment the recorder failed to move the tape altogether (near the low tape sensor). We cannot dismout the tape; it looks as we have a serious problem.

Medicina (Morsiani at Fri Jun 5 13:42:38 METDST 1998) Success
It looks that log file is missing.

Onsala (John Conway at Fri Jun 5 14:17:20 METDST 1998) Failure
Did not observe this experiment, had a complete failure of the MAT card used for head positioning during previous expt. This card had now (Friday midday) been replace and we are observing again. Sorry Franco this expt seems jinxed somehow.

Noto (C.Trigilio at Mon Jun 8 17:38:25 METDST 1998) Success

Jodrell (Alastair Gunn at Tue Jun 9 03:02:41 METDST 1998) Success
High signal in VC4, doesnt look like RFI. No other known problems.

Cambridge (Alastair Gunn at Tue Jun 9 03:03:20 METDST 1998) Minor failures
Vacuum lost at 00:58 UT, tape started again at 12R scan at 01:10 UT. Lost again at 02:00 UT, started at 17F at 02:06 UT. Receiver cryostat failure during the previous night - expect low signal. Telescope taken offline at 09:00 UT to fix receiver.

Calibration (Jean\-Francois Desmurs at Wed Jun 10 15:57:57 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name em025.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 14:52:51 METDST 1998) Success

Urumqi (Liu Xiang at Fri Jun 19 06:03:13 METDST 1998) Minor failures
It was cloud, the clock offset was 1.71us. UT: 1:10--1:45, the antenna problem, and then the antenna was pointing the source2147+145, source 2134+00 was not observed from UT: 1:45.